the base of a triangle is 8cm greater than the height. the area is 42 cm spuared. find the height and length of the base.
Answer by mananth(16949)

let height = h
base = h+8
..
Area of traingle = 1/2 * base * height
42= 1/2 * h * (h+8)
42=1/2*(h^2+8h)
84= h^2+8h
h^2+8h-84=0
h^2+14h-6h-84=0
h(h+14)-6(h+14)=0
(h+14)(h-6)=0
h= 6 Or -14
ignore negative value
h=6 cm
base = h+8=14 cm